### v2

# A phase transition in Barak-Erdős random graphs

Abstract
We study monotone paths in Erdős-Rényi random graphs on numbered vertices. This model appeared in [3] where Benjamini & Tzalik established a phase transition at $p=\frac{\log n}{n}$. We refine the critical value to $p=\frac{\log n-\log\log n}{n}$ and identify the critical window of order $\Theta(1/n)$.
1 Introduction
Recall the definition of a Barak-Erdős random graph $G_{p}$ on the positive integers $\mathbb{N}$: A directed edge $(i,j)$ between any two vertices $i<j\in\mathbb{N}$ exists independently with probability $p\in(0,1)$. We denote this law by $\mathbb{P}_{p}$. We say that there is a monotone path from $1$ to $n$, write $1\nearrow n$, if there exists a path $1=i_{1}\to i_{2}\to\cdots\to i_{k}=n$ in $G_{p}$ whith $i_{1}<i_{2}<\cdots<i_{k}$. We establish a phase transition for the increasing event
$\left\{1\nearrow n\textnormal{ in }G_{p}\right\}.$
Theorem 1.1 (Critical window).
The critical window of the event $\left\{1\nearrow n\textnormal{ in }G_{p_{n}}\right\}$ is of order $\Theta(1/n)$ around $\frac{\log n-\log\log n}{n}$. More precisely, if $x\in\mathbb{R}$
 $\mbox{ for }p_{n,x}=\frac{\log n-\log\log n+x}{n}\quad\mbox{ we have }\quad\mathbb{P}_{p_{n,x}}\left(1\nearrow n\right)\xrightarrow[n\to\infty]{}1-e^{e^{-x}-x}\int_{e^{-x}}^{\infty}\frac{e^{-t}}{t}\mathrm{d}t.$
Remark 1.2.
An alternative expression is
$\mathbb{P}_{p_{n,x}}\left(1\nearrow n\right)\xrightarrow[n\to\infty]{}\frac{1}{2}+\frac{1}{2}\mathbb{E}\left(\tanh\frac{x-\mathfrak{G}}{2}\right)$
for a standard Gumbel variable $\mathfrak{G}$ with density $p_{\mathfrak{G}}(z)=\exp(-(z+e^{-z}))$.
The location and width of the phase transition might be guessed using the expected number of monotone paths in $G_{p}$. Indeed, a simple counting exercise shows that
$\mathbb{E}_{p}\left(\#\textnormal{ paths }1\nearrow n\right)=\sum_{k=0}^{n-2}\binom{n-2}{k}p^{k+1}=p(1+p)^{n-2}$
and for $p\equiv p_{n,x}$ the above display converges to $e^{x}$ as $n\to\infty$.
[…]
Figure 1: A segment from a sampled graph $G_{p}$ satisfying $1\nearrow 4$ and $1\nearrow 8$.
Erdős-Rényi graphs on numbered vertices $\{1,\ldots,n\}$ appeared as an ingredient to recover point configurations in [3]. They prove a sharp threshold at $p=\frac{\log n}{n}$ for the event $1\nearrow n$. Directed acyclic Erdös-Rényi graphs were introduced in [2] and studied as Barak-Erdős graphs since. The articles [8], [9] and [7] established maximal path lengths in Barak-Erdős graphs, both for fixed $p\in(0,1)$ and $p\to 0$.
Numbered vertices can be interpreted as time stamps: each vertex becomes available from its assigned time onward. Related models include random temporal graphs, where monotonicity of paths is considered along edges instead of vertices, see for instance [1], [5] and [4]. These works focus on the length of monotone paths and identify several transition phenomena around $p=c\frac{\log n}{n}$ for both longest and shortest such paths.
A vertex order appears naturally if the nodes correspond to events unfolding in time, as for example in graphical models of causality [10]. Erdős-Rényi graphs directed according to a random vertex order are of particular interest to research in causal discovery [11], where they are widely used in the benchmarking of algorithms that learn causal graphs from data.
Acknowledgments
The first and fourth authors are supported by the ANR grant BISCOTTE - ANR-19-CHIA-0021. The second and third authors are supported by the SuPerGRandMa, the ERG Consolidator no 101087572.
2 Proof
The central ingredient to prove Theorem 1.1 is the introduction of an exploration process. It indicates the position of having passed a certain number of vertices that are reachable from $1$ by a monotone path. We show that, after suitable centering and scaling, this exploration process converges to a Gumbel variable with a deterministic drift in Section 2.1. This allows to compute explicitly the probability that a monotone path $1\nearrow n$ exists.
Define
$N_{k}\coloneq\#\left\{1\leq i\leq k:1\nearrow i\right\}$
the number of vertices up to $k$ that are reachable from $1$ by a monotone path and set $\mathcal{F}_{k}=\sigma(N_{1},\ldots,N_{k})$. A new vertex $k$ is reachable from $1$ precisely when there is an edge to at least one of the $N_{k-1}$ previously reachable vertices. Hence
$\mathbb{P}_{p}\left(N_{k}=N_{k-1}+1|\mathcal{F}_{k-1}\right)=1-\mathbb{P}_{p}\left(N_{k}=N_{k-1}|\mathcal{F}_{k-1}\right)=1-(1-p)^{N_{k-1}}$
and
$\mathbb{P}_{p}(1\nearrow k)=\mathbb{E}\left(1-(1-p)^{N_{k-1}}\right).$ (2.1)
The probability that successive gaps between reachable vertices have sizes $k_{1},\ldots,k_{l}\geq 1$ is
$\mathbb{P}_{p}\!\left(\overset{1}{\bullet}\,\underbrace{\cdots}_{k_{1}}\,\bullet\,\underbrace{\cdots}_{k_{2}}\,\bullet\,\cdots\cdots\,\bullet\,\underbrace{\cdots}_{k_{l}}\,\bullet\right)=\prod_{i=1}^{l}((1-p)^{i})^{k_{i}}\left(1-(1-p)^{i}\right)$
where each large bullet marks a vertex reachable from $1$ and each group of small dots represents the vertices in the corresponding gap: The $i$-th gap is a geometric variable with parameter $1-(1-p)^{i}$, and the gaps are pairwise independent.
2.1 Exploration process
Fix $p\in(0,1)$ and consider the monotone exploration process $(P_{p}(a))_{a>0}$ where
$P_{p}(a)\coloneq\inf\left\{k\geq 1:N_{k}=\lfloor a/p\rfloor\right\}$
denotes the position of the $\lfloor a/p\rfloor$-th vertex reachable from $1$ by a monotone path. Since $P_{p}(a)$ is obtained by summing the successive gap lengths, we have
$P_{p}(a)=\sum_{i=1}^{\lfloor a/p\rfloor}X_{i},\quad X_{i}\sim\mathrm{Geom}\left(1-(1-p)^{i}\right)\text{ independent}.$
The exploration process $(P_{p}(a))_{a>0}$ centered and rescaled, converges, as $p\to 0$, to a standard Gumbel variable $\mathfrak{G}$ with a deterministic drift depending on $a$.
Proposition 2.1 (Convergence of the monotone exploration process).
We have
$\left(pP_{p}(a)-\log\frac{1}{p}\right)_{a>0}\xrightarrow[p\to 0]{(d)}\left(\mathfrak{G}+\log(e^{a}-1)\right)_{a>0}$
for the uniform convergence over every compact subset of $(0,\infty)$.
Proof of Proposition 2.1.
Define $q_{i}\coloneq 1-(1-p)^{i}$. Assume without loss of generality that $p<1/2$. For any $p<1/2$ and $i\leq\lfloor a/p\rfloor$ it holds $q_{i}\leq 1-(1-p)^{\lfloor a/p\rfloor}\leq B(a)<1$ for some constant $B(a)$ independent of $p$. Each $P_{p}(a)$ is the sum of independent geometric variables with expectation $1/q_{i}$.
We first approximate the rescaled geometric variables $q_{i}X_{i}$ by i.i.d. exponential variables via a coupling: Let $(U_{i})_{i}$ be i.i.d. uniform variables on $[0,1]$. Set
$Y_{i}=-\log U_{i}\,\textnormal{ and }\,X_{i}=\inf\left\{k\geq 1:U_{i}\geq(1-q_{i})^{k}\right\}$
so that $Y_{i}$ are exponential variables with intensity $1$ and $X_{i}$ are geometric variables with mean $1/q_{i}$. Then $\log\left(\frac{1}{1-q_{i}}\right)(X_{i}-1)\leq Y_{i}$ which allows the bound
$\left|q_{i}X_{i}-Y_{i}\right|\leq\left|\frac{q_{i}}{\log(1/(1-q_{i}))}-1\right|Y_{i}+q_{i}\leq C(a)q_{i}(Y_{i}+1)$ (2.2)
for some constant $C(a)>0$, using that $q_{i}<B(a)<1$.
In a second step, we express the exploration process using the exponential variables $Y_{i}$ and show that the occurring error has converging expectation and vanishing variance, that is
$pP_{p}(a)-\log\frac{1}{p}$ $\overset{(d)}{=}\sum_{i=1}^{\lfloor a/p\rfloor}\frac{1}{i}Y_{i}-\log\frac{1}{p}+R_{p}(a)$
with rest term
$R_{p}(a)=\sum_{i=1}^{\lfloor a/p\rfloor}p\left(\frac{1}{q_{i}}-\frac{1}{ip}\right)Y_{i}+\sum_{i=1}^{\lfloor a/p\rfloor}\frac{p}{q_{i}}(q_{i}X_{i}-Y_{i}).$
Lemma 2.2.
As $p\to 0$, the rest term satisfies
$\mathbb{E}(R_{p}(a))=\log(e^{a}-1)-\log a+o(1)\,\text{ and }\,\operatorname{Var}(R_{p}(a))=o(1).$
In a last step, we use the classical fact that the sum of the rescaled exponential variables, properly centered, converges in law to a standard Gumbel variable,
$\sum_{i=1}^{\lfloor a/p\rfloor}\frac{1}{i}Y_{i}-\log\frac{1}{p}\xrightarrow[p\to 0]{(d)}\mathfrak{G}+\log a,$ (2.3)
see e.g. [6, Lemma 11.2]. Therefore
$pP_{p}(a)-\log\frac{1}{p}$ $\xrightarrow[p\to 0]{(d)}\mathfrak{G}+\log(e^{a}-1)$
for any fixed $a>0$. In fact, this is the same Gumbel variable for any two $a_{1},a_{2}>0$.
We now conclude the convergence as a process in $a>0$ via a probabilistic version of Dini’s lemma, see [6, Theorem 11.6]: The process
$\left(pP_{p}(a)-\log\frac{1}{p}-\mathfrak{G}\right)_{a>0}$
is increasing in $a$ and for every fixed $a>0$ it converges in law to $\log(e^{a}-1)$ which is an increasing continuous function. Therefore Dini’s lemma implies the convergence as a process
$\left(pP_{p}(a)-\log\frac{1}{p}-\mathfrak{G}\right)_{a>0}\xrightarrow[p\to 0]{(d)}\left(\log(e^{a}-1)\right)_{a>0}$
for the topology of uniform convergence over every compact subset of $(0,\infty).$
It remains to prove Lemma 2.2.
Proof of Lemma 2.2.
For the expectation notice that $\mathbb{E}(q_{i}X_{i})=1=\mathbb{E}(Y_{i})$ so that
$\mathbb{E}(R_{p}(a))$ $=\mathbb{E}\left(\sum_{i=1}^{\lfloor a/p\rfloor}p\left(\frac{1}{q_{i}}-\frac{1}{ip}\right)Y_{i}\right)=p\sum_{i=1}^{\lfloor a/p\rfloor}\frac{1}{q_{i}}-\frac{1}{ip}$
$=p\sum_{i=1}^{\lfloor a/p\rfloor}\left(\frac{1}{1-e^{-ip}}-\frac{1}{ip}\right)+p\sum_{i=1}^{\lfloor a/p\rfloor}\left(\frac{1}{q_{i}}-\frac{1}{1-e^{-ip}}\right)$
$\leq\int_{0}^{a}\left(\frac{1}{1-e^{-s}}-\frac{1}{s}\right)\mathrm{d}s+p\sum_{i=1}^{\lfloor a/p\rfloor}\frac{C(a)}{i}$
$=\log(e^{a}-1)-\log a+o(1)$
for some constant $C(a)>0$ depending on $a$. For the variance write
$Z_{i}=\left(\frac{1}{q_{i}}-\frac{1}{ip}\right)Y_{i}+\frac{1}{q_{i}}(q_{i}X_{i}-Y_{i})\leq c_{1}(a)Y_{i}+c_{2}(a)$
for some constants $c_{1}(a)>0$ and $c_{2}(a)>0$, using the bound $\left(\frac{1}{q_{i}}-\frac{1}{ip}\right)\leq\frac{a}{2(1-e^{-a})}$ for $i\leq\lfloor a/p\rfloor$ and (2.2). Then, as $Y_{i}\sim\operatorname{Exp}(1)$ for all $i$, this yields
$\operatorname{Var}(Z_{i})\leq\mathbb{E}\left(Z_{i}^{2}\right)=\mathcal{O}(1).$
Independence of the $Z_{i}$ thus yields
$\operatorname{Var}\left(R_{p}(a)\right)=p^{2}\sum_{i=1}^{\lfloor a/p\rfloor}\operatorname{Var}(Z_{i})=o(1).$
2.2 Critical window
We now choose $p$ dependent on $n$ and let $n\to\infty$. More precisely, let $b>0$ and let $p=p_{n}$ satisfy
$n=\frac{\log(b/p_{n})}{p_{n}}.$ (2.4)
Proof of Theorem 1.1.
For each $n\in\mathbb{N}$ define the random variable $\mathfrak{X}_{n}$ as the number of vertices smaller than $n$ that are reachable from $1$ by a monotone path, scaled by $p_{n}$, that is
$P_{p_{n}}(\mathfrak{X}_{n})=n\quad\text{with}\quad\lfloor\mathfrak{X}_{n}/p_{n}\rfloor=N_{n}.$
Then, by (2.1)
$\mathbb{P}_{p_{n}}\left(1\nearrow n\right)=\mathbb{E}\left(1-(1-p_{n})^{\lfloor\mathfrak{X}_{n}/p_{n}\rfloor-1}\right)=\mathbb{E}\left(1-e^{-\mathfrak{X}_{n}}(1+o(1))\right).$ (2.5)
On the other hand, by the convergence in law from Proposition 2.1 and the Skorokhod representation theorem there is a probability space with a variable $\tilde{\mathfrak{G}}$ and a process $\tilde{P}_{p_{n}}$ inducing variables $\tilde{\mathfrak{X}}_{n}$ with the same law as $\mathfrak{G}$, $P_{p_{n}}$ and $\mathfrak{X}_{n}$ respectively such that almost-sure-convergence
$np_{n}-\log\frac{1}{p_{n}}-\log\left(e^{\tilde{\mathfrak{X}}_{n}}-1\right)\xrightarrow[n\to\infty]{}\tilde{\mathfrak{G}}\quad\text{a.s.}$
holds. Therefore
$1-e^{-\tilde{\mathfrak{X}}_{n}}\xrightarrow[n\to\infty]{}1-\frac{1}{1+be^{-\tilde{\mathfrak{G}}}}\quad\text{a.s.}$ (2.6)
and this convergence holds in law for $\mathfrak{X}$ and $\mathfrak{G}$. Combine (2.5) and (2.6) to compute
$\mathbb{P}_{p_{n}}\left(1\nearrow n\right)=1-\mathbb{E}\left(\frac{1}{1+be^{-\mathfrak{G}}}\right)+o(1)=1-\frac{1}{b}e^{1/b}\int_{1/b}^{\infty}\frac{e^{-t}}{t}\mathrm{d}t+o(1).$ (2.7)
The function $f(b)=\frac{1}{b}e^{1/b}\int_{1/b}^{\infty}\frac{1}{t}e^{-t}\mathrm{d}t$ is continuous on $(0,\infty)$ with asymptotics
$f(b)\xrightarrow{b\to\infty}0\quad\text{and}\quad f(b)\xrightarrow{b\to 0}1.$
So for any $b>0$ the choice $p_{n}$ in (2.4) yields a non-trivial probability for $\{1\nearrow n\}$. It remains to determine the matching value $b=b_{x}$ when choosing $p_{n,x}$ as in the theorem. From (2.4) we have
$b_{x}=p_{n,x}e^{np_{n,x}}=\left(\frac{\log n-\log\log n}{n}+\frac{x}{n}\right)\frac{n}{\log n}e^{x}=\left(1+\frac{x-\log\log n}{\log n}\right)e^{x}\sim e^{x}.$
Plug this into equation (2.7). ∎
Remark 2.3 (Number of paths).
Denote by $M_{a}^{+}$ the number of paths $1\nearrow P_{p}(a)$, conditioned to be positive. We conjecture that, as $p\to 0$, $M_{a}^{+}$ converges to a geometric random variable with parameter $e^{-a}$. Heuristically, it satisfies the recursive relation
$M_{a}^{+}\overset{(d)}{=}\sum_{i=1}^{W}M_{aU_{i}}^{+}$
where $W$ is a $\mathrm{Poisson}(a)$ random variable conditioned to be positive, and $(U_{i})_{i\geq 1}$ are i.i.d $\mathrm{Unif}[0,1]$ variables.